OpenStudy (anonymous):

80% is equivalent to a.) 8.0 and 8/10 b.) 0.8 and 10/8 c.) 0.8 and 8/10

OpenStudy (anonymous):

oh, is it 0.8 and 8/10?

OpenStudy (mathstudent55):

% means \( \times \dfrac{1}{100} \) Can you multiply 80 and \(\dfrac{1}{100} \) together? What do you get?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Sweet. Thanks!

OpenStudy (mathstudent55):

\(80\% = 80 \times \dfrac{1}{100} = \dfrac{80}{1} \times \dfrac{1}{100} = \dfrac{80}{100} = \color{red}{\dfrac{8}{10} = 0.8}\)

OpenStudy (mathstudent55):

wlcm